OpenStudy (anonymous):

x^2-5x-24=0 quadratic equation...explain while solving? step by step?

OpenStudy (anonymous):

factorizing: x^2-5x-24 = (x-8)(x+3) = 0 so x=8 and x=-3 are the roots

OpenStudy (anonymous):

how?how do you get x-8 and x+3?explain?

OpenStudy (anonymous):

factorizing..... if you multiplie the right side you get the same what's on the left. If you notice: (x-a)(x-b) = x2 -bx -ax +ab = x2 - (a+b)x +ab. So the sum with oposite sign is the x coeficient and the free term is the product. So just check posible factors of 24...

OpenStudy (anonymous):

You just posted this question and two of us helped you here: http://openstudy.com/study#/updates/4f8d707be4b00280a9c36914

OpenStudy (anonymous):

i know GT,but simply giving the answer doesn't always suffice..

OpenStudy (anonymous):

Then, you are supposed to ask what part of it you did not get. You never said anything.

OpenStudy (anonymous):

@GT is right. He gave you same explanation like i did....

OpenStudy (anonymous):

if you read the question,it says "explain while solving",not simply give the answer.

OpenStudy (anonymous):

Rohangrr did a fine job too. No response to either of us. Re-post the same question.

OpenStudy (anonymous):

@Rohangrr I meant.

OpenStudy (anonymous):